# Quantum entanglement in phase space

### 09 Mar 2026 v3

@@ Introduction
where $\hat{x}_{k}$ and $\hat{p}_{k}$ are respectively the position and momentum quadratures of system $k\in\{A,B\}$ satisfying $\hat{x}_{k},\hat{p}_{l}]=2i\delta_{kl}$. Observing a violation of Eq. (1) thus implies entanglement between $A$ and $B$. Moreover, a simple generalization of this condition has been shown to be necessary and sufficient for detecting entanglement in Gaussian states [2].
@@ Results
Wigner function formalism.– Consider a system of $N$ bosonic modes, each described by creation and annihilation operators $\hat{a}_{k}^{\dagger}$ and $\hat{a}_{k}$, which satisfy the commutation relation $\left[\hat{a}_{k},\hat{a}_{k^{\prime}}^{\dagger}\right]=\delta_{kk^{\prime}}$ for $k,k^{\prime}\in\{1,\dots,N\}$. For each mode, it is natural to introduce position and momentum operators $\hat{x}_{k}=\hat{a}_{k}^{\dagger}+\hat{a}_{k}$ and $\hat{p}_{k}=i\left(\hat{a}_{k}^{\dagger}-\hat{a}_{k}\right)$, satisfying $\hat{x}_{k},\hat{p}_{k^{\prime}}]=2i\delta_{kk^{\prime}}$. We arrange these in the vector $\hat{\bm{\xi}}=\left(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N}\right)^{\top}$, which is associated to the vector of phase space coordinates $\bm{\xi}\equiv\left(x_{1},p_{1},\ldots,x_{N},p_{N}\right)^{\top}$. A convenient representation of the state $\rho$ of a CV system is given by the Wigner function
@@
The rest of this work will deal with a bipartite, i.e. $N=2$, system described by the density matrix $\rho_{AB}$ or, equivalently, by the joint Wigner function $W_{AB}(x_{A},p_{A},x_{B},p_{B})$. Our goal is to detect entanglement between the $A$ and $B$ subsystems based on this function. Crucially, as illustrated in Fig. 1, our criteria do not require characterizing the entire four-dimensional Wigner function. Instead, they rely on specific two-dimensional slices where the variables $(x_{B},p_{B})$ are constrained to be functions of $(x_{A},p_{A})$, reducing the measurement complexity from four independent variables to just two. (For notational simplicity, we then drop the subscript $A$.) Note that such a result is also useful towards revealing multipartite entanglement, as one can test inseparability across all possible bipartitions of the $N>2$ subsystems.
@@
where in the last row we substituted $x\rightarrow x+X\tan\theta$ and similarly for $p$. This can be interpreted as follows: Perform a beam-splitter transformation $\left(\begin{smallmatrix}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{smallmatrix}\right)$ to the two modes $A$ and $h(B)$, and denote the resulting modes as $A^{\prime}:(x,p)$ and $B^{\prime}:(X,P)$. Notice that Eq. (11) is exactly the Wigner function of mode $B^{\prime}$, with mode $A^{\prime}$ integrated out. Therefore Eq. (11) should be upper-bounded by $1/(2\pi)$ [25]. Detailed proof can be found in Sec. IV of SM [22]. $\Box$
@@ III. An intuitive phase-space counterpart to the Duan-Simon criterion and their comparison
Using Williamson’s theorem, the covariance matrix of a single-mode Gaussian state can be diagonalized as $\sigma=\operatorname{diag}(l,l)$ through a symplectic transformation, where $l\geqslant 1$ due to the uncertainty relation. Conversely, if $0<l<1$ the purity $1/\sqrt{\operatorname{Det}\sigma}$ would exceed $1$. Consequently, from the purity of the partially transposed state we can establish an entanglement criterion based on the phase space distribution.
@@
Proof. If $\rho_{AB}$ is separable, then its partial transpose $\rho_{AB}^{T_{B}}$ remains a physical state [1]. Applying a beam-splitter transformation to $\rho_{AB}^{T_{B}}$ must also result in a physical state with its purity being upper bounded by $1$. ∎
@@
Without losing generality, assume $n>m$. The purity reaches its maximum value of $2/(m+n-\sqrt{4c^{2}+(n-m)^{2}})$ when choosing $\tan(2\theta)=2c/(n-m)$. The value exceeds $1$ when $c^{2}>(n-1)(m-1)$, indicating that $\rho_{G}$ is entangled. Thus, the conditions derived from both methods are equivalent.
@@ V. Criterion (I) is sufficient and necessary for Gaussian states
The value of $I$ reaches its maximum
@@ VIII. Detailed analysis of some exemplary quantum states
where, $N$ is the number of modes, $\vec{x}=(x_{1},p_{1},\cdots,x_{N},p_{N})^{T}$ is a phase-space coordinate vector, and $\vec{\xi}$ is the displacement vector. We only consider the two-mode case ($N=2$) in our work. The analytical expressions of Wigner function for TMST states is obtained
@@ B. How to measure the criteria
Each individual displaced-parity measurement yields as outcome ±1, and averaging over many measurement repetitions provides the value of the Wigner function at the corresponding point in phase-space. For a total finite number of measurements $N$, let each measurement result be a random variable $X_{i}\in\{+1,-1\}$ with mean value $E[X_{i}]=\mu$ (which is proportional to $W(\alpha;\beta)$). The sample mean is

### 15 Mar 2026 v4

@@ Introduction
where $\hat{x}_{k}$ and $\hat{p}_{k}$ are respectively the position and momentum quadratures of system $k\in\{A,B\}$ satisfying $\hat{x}_{k},\hat{p}_{l}]=2i\delta_{kl}$. Observing a violation of Eq. (1) thus implies entanglement between AA and $B$. Moreover, a simple generalization of this condition has been shown to be necessary and sufficient for detecting entanglement in Gaussian states [2].
@@ Results
Wigner function formalism.– Consider a system of NN bosonic modes, each described by creation and annihilation operators $\hat{a}_{k}^{\dagger}$ and $\hat{a}_{k}$, which satisfy the commutation relation $\left[\hat{a}_{k},\hat{a}_{k^{\prime}}^{\dagger}\right]=\delta_{kk^{\prime}}$ for $k,k^{\prime}\in\{1,\dots,N\}$. For each mode, it is natural to introduce position and momentum operators $\hat{x}_{k}=\hat{a}_{k}^{\dagger}+\hat{a}_{k}$ and $\hat{p}_{k}=i\left(\hat{a}_{k}^{\dagger}-\hat{a}_{k}\right)$, satisfying $\hat{x}_{k},\hat{p}_{k^{\prime}}]=2i\delta_{kk^{\prime}}$. We arrange these in the vector $\hat{\bm{\xi}}=\left(\hat{x}_{1},\hat{p}_{1},\ldots,\hat{x}_{N},\hat{p}_{N}\right)^{\top}$, which is associated to the vector of phase space coordinates $\bm{\xi}\equiv\left(x_{1},p_{1},\ldots,x_{N},p_{N}\right)^{\top}$. A convenient representation of the state $\rho$ of a CV system is given by the Wigner function
@@
The rest of this work will deal with a bipartite, i.e. $N=2$, system described by the density matrix $\rho_{AB}$ or, equivalently, by the joint Wigner function $W_{AB}(x_{A},p_{A},x_{B},p_{B})$. Our goal is to detect entanglement between the AA and $B$ subsystems based on this function. Crucially, as illustrated in Fig. 1, our criteria do not require characterizing the entire four-dimensional Wigner function. Instead, they rely on specific two-dimensional slices where the variables $(x_{B},p_{B})$ are constrained to be functions of $(x_{A},p_{A})$, reducing the measurement complexity from four independent variables to just two. (For notational simplicity, we then drop the subscript AA.) Note that such a result is also useful towards revealing multipartite entanglement, as one can test inseparability across all possible bipartitions of the $N>2$ subsystems.
@@
where in the last row we substituted $x\rightarrow x+X\tan\theta$ and similarly for pp. This can be interpreted as follows: Perform a beam-splitter transformation $\left(\begin{smallmatrix}\cos\theta&\sin\theta\\ \sin\theta&-\cos\theta\end{smallmatrix}\right)$ to the two modes AA and $h(B)$, and denote the resulting modes as $A^{\prime}:(x,p)$ and $B^{\prime}:(X,P)$. Notice that Eq. (11) is exactly the Wigner function of mode $B^{\prime}$, with mode $A^{\prime}$ integrated out. Therefore Eq. (11) should be upper-bounded by $1/(2\pi)$ [25]. Detailed proof can be found in Sec. IV of SM [22]. $\Box$
@@ III. An intuitive phase-space counterpart to the Duan-Simon criterion and their comparison
Using Williamson’s theorem, the covariance matrix of a single-mode Gaussian state can be diagonalized as $\sigma=\operatorname{diag}(l,l)$ through a symplectic transformation, where $l\geqslant 1$ due to the uncertainty relation. Conversely, if $0<l<1$ the purity $1/\sqrt{\operatorname{Det}\sigma}$ would exceed 11. Consequently, from the purity of the partially transposed state we can establish an entanglement criterion based on the phase space distribution.
@@
Proof. If $\rho_{AB}$ is separable, then its partial transpose $\rho_{AB}^{T_{B}}$ remains a physical state [1]. Applying a beam-splitter transformation to $\rho_{AB}^{T_{B}}$ must also result in a physical state with its purity being upper bounded by 11. ∎
@@
Without losing generality, assume $n>m$. The purity reaches its maximum value of $2/(m+n-\sqrt{4c^{2}+(n-m)^{2}})$ when choosing $\tan(2\theta)=2c/(n-m)$. The value exceeds 11 when $c^{2}>(n-1)(m-1)$, indicating that $\rho_{G}$ is entangled. Thus, the conditions derived from both methods are equivalent.
@@ V. Criterion (I) is sufficient and necessary for Gaussian states
The value of II reaches its maximum
@@ VIII. Detailed analysis of some exemplary quantum states
where, NN is the number of modes, $\vec{x}=(x_{1},p_{1},\cdots,x_{N},p_{N})^{T}$ is a phase-space coordinate vector, and $\vec{\xi}$ is the displacement vector. We only consider the two-mode case ($N=2$) in our work. The analytical expressions of Wigner function for TMST states is obtained
@@ B. How to measure the criteria
Each individual displaced-parity measurement yields as outcome ±1, and averaging over many measurement repetitions provides the value of the Wigner function at the corresponding point in phase-space. For a total finite number of measurements NN, let each measurement result be a random variable $X_{i}\in\{+1,-1\}$ with mean value $E[X_{i}]=\mu$ (which is proportional to $W(\alpha;\beta)$). The sample mean is